2. At the beginning of the year, BIBI Company’s liabilities equal P200,000. During the year, assets increase by P80,000, and at year-end assets equal P460,000. Liabilities decrease P50,000 during the year. What are the beginning and ending amounts of equity?


How much is the beginning capital?


1
Expert's answer
2021-10-05T11:54:35-0400

If assets increase by 40000 and liabilities decrease by 30000 then it means that net assets increased by 70000 (i.e. 40000 - (-30000) ). Accordingly, equity increased by 70000

So equity = 96000 + 70000 = P166,000

This can be rechecked from the accounting equation i.e.

Assets = Equity + Liabilties

Initial equation

Assets = 96000 + 100000

196000 = 96000 + 100000

Increasing Assets by 40K and decreasing Liabilities by 30K

(196000 + 40000) = Equity + (100000–30000)

Solving this we get:

Equity = P166,000